the sum of the digit of a two digit number is 11 the digit at one's place is 2 more than twice the digit in tens place then find the number
Answer by lwsshak3(11628)

the sum of the digit of a two digit number is 11 the digit at one's place is 2 more than twice the digit in tens place then find the number
**
let u=units digit
let t=tens digit
u+t=11
t=11-u
..
u=2t+2
u=2(11-u)+2
u=22-2u+2
3u=24
u=8
t=11-u=3
number: 10t+u=30+8=38